OpenStudy (anonymous):

If f(x) = 2x+1 and g(x) = 3x+4, what is f(g(4))?

OpenStudy (anonymous):

33

OpenStudy (anonymous):

3*4+4=16 2*16+1=33

OpenStudy (anonymous):

Do what is inside the () first..so g(4) = 3(4) + 4 = 16 now do F(16) which is 2(16) + 1 = 33